OpenStudy (anonymous):

Owen scored 85 points on a 100-point test and 26 on a 30-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

85 points on a 100 point test = 85/100 = 85% 26 points on a 30 point test = 26/30 = _%?